# A Common Origin of Asymmetric Self-interacting Dark Matter and Dirac Leptogenesis

### v1

@@ Appendix B DM Self-interaction Cross-sections at Low Energy
$\sigma^{\rm classical}_{T}(\rm Attractive)=\left\{\begin{array}[]{l}\frac{4\pi}{M^{2}_{Z^{\prime}}}\beta^{2}\ln(1+\beta^{-1});~{}~{}\beta\leqslant 10^{-1}\\ \frac{8\pi}{M^{2}_{Z^{\prime}}}\beta^{2}/(1+1.5\beta^{1.65});10^{-1}\leq\beta\leqslant 10^{3}\\ \frac{\pi}{M^{2}_{Z^{\prime}}}(\ln\beta+1-\frac{1}{2}ln^{-1}\beta);~{}~{}\beta\geq 10^{3}\\ \end{array}\right.$ (24)
@@
$\sigma^{\rm classical}_{T}(\rm Repulsive)=\left\{\begin{array}[]{l}\frac{2\pi}{M^{2}_{Z^{\prime}}}\beta^{2}\ln(1+\beta^{-2});~{}~{}\beta\leqslant 1\\ \frac{\pi}{M^{2}_{Z^{\prime}}}(\ln 2\beta^{2}-\ln~{}\ln 2\beta)^{2};~{}~{}\beta\geq 1\\ \end{array}\right.$ (25)
@@
$\lambda_{\pm}=\left\{\begin{array}[]{l}1+\frac{iM_{\chi}v}{2kM_{Z^{\prime}}}\pm\sqrt{\frac{\alpha_{D}M_{\chi}}{kM_{Z^{\prime}}}-\frac{M^{2}_{\chi}v^{2}}{4k^{2}M^{2}_{Z^{\prime}}}}~{}~{}~{}~{}{\rm Attractive}\\ 1+\frac{iM_{\chi}v}{2kM_{Z^{\prime}}}\pm i\sqrt{\frac{\alpha_{D}M_{\chi}}{kM_{Z^{\prime}}}+\frac{M^{2}_{\chi}v^{2}}{4k^{2}M^{2}_{Z^{\prime}}}}~{}~{}~{}~{}{\rm Repulsive}\\ \end{array}\right.$ (28)

### v2

@@ Appendix B DM Self-interaction Cross-sections at Low Energy
[…]
@@
[…]
@@
[…]
@@
Appendix C Washout Analysis
In this appendix we provide a detailed analysis of why the three-body decay $\psi\to\chi f\bar{f}$ and portal interactions do not wash out either the dark-sector or visible-sector asymmetries.
After the heavy scalar doublets $\xi_{i}$ have decayed and generated the primordial $L$ and $D$ asymmetries, the remaining dark-sector dynamics involves the following processes:
  • the three-body decay and inverse decay
    $\psi\;\leftrightarrow\;\chi f\bar{f},$
  • $Z^{\prime}$-mediated dark–visible scatterings
    $\psi f\leftrightarrow\chi f,\qquad\chi\chi\leftrightarrow f\bar{f},\qquad\psi\bar{\psi}\leftrightarrow f\bar{f},$
  • the Yukawa portal interaction $\lambda\,\psi\,\eta\,\chi$, which is irrelevant once $\eta$ becomes Boltzmann suppressed.
For washout to occur, two conditions must be satisfied: (i) the relevant interactions must violate the quantum number that carries the asymmetry (here the dark charge $D$ or the lepton number $L$), and (ii) the violating processes must be in chemical equilibrium, $\Gamma\gtrsim H$, so that the corresponding chemical potentials are driven to zero. As shown below, neither condition is realised in our model after the decay epoch of the heavy scalar doublets.
Three body decay and inverse decay $\chi f\bar{f}\leftrightarrow\psi$
The three-body decay $\psi\to\chi f\bar{f}$ is induced by the soft $\mathcal{Z}_{2}$-breaking operator $\mu^{2}\,\eta^{\dagger}H$. Crucially, although this term breaks the discrete parity, it preserves the dark charge $D$:
$D(\psi)=D(\chi),\qquad D(f)=D(\bar{f})=0.$
Therefore the decay merely transfers the primordial $\psi$ asymmetry into the stable state $\chi$, and the total
$Y_{\Delta D}=Y_{\Delta\psi}+Y_{\Delta\chi}$
is conserved exactly.
Inverse decays $\chi f\bar{f}\to\psi$ are also $D$-conserving and hence cannot erase $Y_{\Delta D}$. Furthermore, even ignoring charge considerations, they are extremely suppressed. Using the three-body decay width,
$\Gamma_{\psi}\sim\frac{y_{\rm eff}^{4}}{192\pi^{3}}\,\frac{M_{\psi}^{5}}{\Lambda^{4}},$
we get the scaling relation,
$\boxed{\Gamma_{\Psi}\;\simeq\;1.68\times 10^{-29}\ {\rm GeV}\;y_{\rm eff}^{4}\,\left(\frac{M_{\psi}}{1~{\rm TeV}}\right)^{5}\left(\frac{10^{10}~{\rm GeV}}{\Lambda}\right)^{4}}$
Even for the most conservative values,
$y_{\rm eff}=1,\quad M_{\psi}=1~\mathrm{TeV},\quad\Lambda=10^{10}~\mathrm{GeV},$
one obtains
$\Gamma_{\psi}\simeq 1.68\times 10^{-29}~\mathrm{GeV},\qquad\frac{\Gamma_{\psi}}{H(T=M_{\psi})}\simeq 10^{-13}.$
The inverse decay rate is further suppressed by SM thermal number densities in a three-body initial state. Thus $\Gamma_{\rm inv}/H\ll 10^{-13}$ and inverse decays are always out of equilibrium.
Portal-mediated scatterings:
The only remaining processes that could communicate between the dark and visible sectors are the $Z^{\prime}$-mediated scatterings,
$\psi f\leftrightarrow\chi f,\qquad\chi\chi\leftrightarrow f\bar{f},\qquad\psi\bar{\psi}\leftrightarrow f\bar{f}.$
These preserve $D$, but if they were fast enough, they could in principle transfer the dark asymmetry into the visible sector and wash out the lepton asymmetry by coupling efficiently to the $L$-violating interactions in the $\xi_{i}$ sector.
In our model the $Z^{\prime}$ is a light mediator in the MeV range, and the dark gauge coupling $g^{\prime}$ is relatively large ($g^{\prime}\sim 0.1$) in order to generate the desired self-interacting dark matter cross section. However, the potentially dangerous washout processes are not controlled by $g^{\prime}$ alone, but by the portal coupling of $Z^{\prime}$ to the SM. This portal is very small in our setup from the requirement of direct search. The kinetic mixing parameter $\epsilon$ is required to be very small ($\epsilon\lesssim 10^{-8}$), as shown in Fig. 14 of the paper. As a result, the effective coupling that enters SM-dark scatterings is
$g^{\prime}_{\rm eff}\equiv\epsilon\,g^{\prime}\sim 10^{-9},$
which is many orders of magnitude smaller than the dark matter self-interaction coupling. For relativistic $Z^{\prime}$-mediated scatterings the thermally averaged cross section scales as $\langle\sigma v\rangle\sim(g^{\prime}_{\rm eff})^{4}/T^{2}$, while the number density of SM fermions scales as $n(T)\sim T^{3}$. Thus the scattering rate may be written as
$\Gamma_{\rm scatt}(T)=n(T)\langle\sigma v\rangle\simeq c_{\rm scatt}\,(g^{\prime}_{\rm eff})^{4}\,T,$
where $c_{\rm scatt}$ is an $\mathcal{O}(10^{-1})$ coefficient encoding spin sums, thermal averaging and numerical phase-space factors. At the epoch of $\psi$ decay ($T\simeq M_{\psi}$), we obtain
$\frac{\Gamma_{\rm scatt}}{H}\simeq\frac{c_{\rm scatt}\,M_{\rm Pl}}{1.66\sqrt{g_{*}}\,M_{\psi}}\,(g^{\prime}_{\rm eff})^{4}\simeq 1.1\times 10^{16}\,(g^{\prime}_{\rm eff})^{4}\left(\frac{10^{3}~\mathrm{GeV}}{M_{\psi}}\right).$
For $M_{\psi}$ at TeV scale, with $g^{\prime}_{\rm eff}\sim 10^{-9}$ (corresponding to $g^{\prime}\sim 0.1$ and $\epsilon\sim 10^{-8}$), this gives
$\frac{\Gamma_{\rm scatt}}{H}\Big|_{T=M_{\psi}}\sim 10^{-20},$
i.e., the scattering rate is too slow compared to the Hubble rate. Thus the dark and visible sectors never approach chemical equilibrium. Potentially dangerous washout channels being many orders of magnitude below the Hubble rate in the allowed parameter region, the asymmetries produced are safely preserved.
Absence of LNV processes after $\xi_{i}$ decays
It is to be mentioned that, even if the portal interactions were faster, they still would not wash out the lepton asymmetry unless they also communicated with lepton-number violating (LNV) processes. But LNV interactions in our model come solely from the decays and inverse decays of the heavy $\xi_{i}$ fields. Once the temperature drops below $T\ll M_{\xi_{i}}$ these interactions are Boltzmann suppressed and effectively decoupled. Thus there is no operational LNV channel for any dark-visible process to couple to, and the visible-sector lepton asymmetry remains protected. Hence neither the dark-sector asymmetry nor the visible lepton asymmetry is affected by late-time decays or scatterings, and no washout occurs anywhere in the parameter space relevant for successful asymmetry generation.